OpenStudy (anonymous):

Which equation represents the line that passes through (–8, 11) and (4,7/2)?

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

y = x + 6 y = x + 16 y = x – 49 y = x + 71

OpenStudy (anonymous):

Feels like y = x + 6

OpenStudy (anonymous):

Thank You!!

OpenStudy (anonymous):

You are welcomed.